OpenStudy (johnnewman):

Katie invested $1200 at 8.5% interest. She earned $612 interest at the end of y years. What is the value of y? A. 18 B. 12 C. 6 D. 3

OpenStudy (benlindquist):

gues?

OpenStudy (johnnewman):

c

OpenStudy (johnnewman):

1020

OpenStudy (benlindquist):

ok leme see

OpenStudy (johnnewman):

you need to make the % into a .

OpenStudy (benlindquist):

(si)=ptr/100 si = simple interest p=1200 r=8.5 (si)=ptr/100 612=1200*t*8.5/100 t=((612*100)/(1200*8.5))

OpenStudy (johnnewman):

ok

OpenStudy (benlindquist):

get it?

OpenStudy (johnnewman):

ok

OpenStudy (benlindquist):

medal pplease

OpenStudy (johnnewman):

but ther is not 60

OpenStudy (johnnewman):

i know 6 go into 60 10 times

OpenStudy (benlindquist):

lee check me work

OpenStudy (johnnewman):

i know 6 go into 60 10 times

OpenStudy (benlindquist):

uh huh

OpenStudy (anonymous):

\[1200(1.085)^y=1812\] solve for \(y\)